OpenStudy (toxicsugar22):

Two of the vertics of rectangle ABCD are A(-5,2) and B(-2,-4) 1.Find an equation for the line that includes side AB. 2.Find and eqaution for the line that includes side BC

OpenStudy (anonymous):

1. We have two points so we can find the slope=(y2-y1)/(x2-x1)=(-4-2)/(-2-(-5))=-6/3=-2. We pick either point and use point slope form. Picking A: y-2=-2*(x+5) y=-2x-8 2. This is a rectangle so side AB and BC are perpendicular. That means their slopes are the negative reciprocal of each other. so the slope of line BC is -1/-2=1/2. We know point B so we use point slope form again. y+4=1/2*(x+2) y=(1/2)x-3

OpenStudy (anonymous):

does it make sense?

OpenStudy (toxicsugar22):

yes

OpenStudy (toxicsugar22):

thanks